OpenStudy (anonymous):

h(x)=xcot(8sqrt(x))+20. Find the derivative.

OpenStudy (anonymous):

My question is why we use the product rule, instead of the chain rule.

OpenStudy (anonymous):

Because isn't it cotangent of x, which in this case is (8(sqrt(x)))?

OpenStudy (anonymous):

It's not cot times that, because cot means nothing by itself i thought? Must be of something.

OpenStudy (raden):

h(x)=xcot(8sqrt(x))+20 let u = x and v = cot(8sqrt(x)) that's why u have to use the product rule h ' = u'v + uv'

OpenStudy (anonymous):

Oh i see, i was letting u=xcot and v=(8sqrtx)

OpenStudy (raden):

to get v', use the chain rule

OpenStudy (anonymous):

I see now.

OpenStudy (anonymous):

Thanks

OpenStudy (raden):

hehe, welcome :)